The 2018 income statement of Anderson Medical Supply Company reported net sales of $10 million, cost of goods sold of $5.6 million, and net income of $840,000. The following table shows the companys comparative balance sheets for 2018 and 2017: 0182017 $460 540 Assets Cash Accounts receivable Inventory Property, plant, and equipment (net) 860 1,300 1,100 3,2002,920 $5,820 $5,140 Total assets Liabilities and shareholders equity Current liabilities Bonds payable Paid-in capital Retained earnings $1,120 990 1,600 1,600 1,800 1,800 1 300 750 Total liabilities and shareholders equity $5,820$5,140 Required: 1. Calculate Andersons turnover ratios for 2018. (Consider 365 days a year. Round your answers to 2 decimal places.) Inventory turnover ratio Receivables tunover ratio Average collection period Asset turnover ratio times times days times
Answer #1

Inventory turnover ratio = COGS / Average inventory

= 5600 / {(1300+1100)/2} = 4.67 times

Receivables turnover ratio = net Credit sale / Average account receivables

= 10000 / {860+580) / 2} = 13.89 times

Average collection period = 365 / Receivables turnover ratio

= 365 / 13.89 = 26.28 days

Asset turnover ratio = Net Sales / Average fixed assets

= 10000 / {(3200+2920) / 2} = 3.27 days